OpenStudy (anonymous):

Solve the following equation. One point is given showing steps and one point given for the correct solution. x/x+3 + 2/5 = 0

OpenStudy (anonymous):

type out your question using the equation button below for more clarity

OpenStudy (anonymous):

\[\frac {x}{x+3} + \frac 2 5 = 0\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

Yes that is it

OpenStudy (anonymous):

okay. Let me give you another question y + 1 = 0 solve for y